Calculate the amount of 92U235 which must go nuclear fission to produce the same amount of energy as produced by 1 ton of coal.  The heat of combustion of coal is 8X103 cal/gram.  The fission of 92U235 produces 200 Mev energy/fission.  (Use 1 cal = 4.2 J, 1 ton = 106gram.  AV number is 6.023X1023 and 1. 602 X 10-19.  J= 1eV)

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Kiran,

combustion of coal is 8X103 cal/gram

1 ton produces=8X103x106x4.2 J

200 MeV = 200 * 106*1. 602 * 10-19.  J/fission

let y fissions

then y*200 * 106x1. 602 X 10-19.=8X103x106x4.2

       y=10.487*1020 atoms

no of moles = y/6.023X1023

                       =0.00174 moles